nightshade0773:

Mr. Morton bought a new shed with 156 more square feet of storage than his old shed. Mr. Morton's new shed has 300 square feet of storage. Let s represent how many square feet of storage Mr. Morton's old shed had. Which equation models the problem? Solve this equation to find how many square feet of storage Mr. Morton's old shed had. square feet some one please help

Extrinix:

So you are looking for the `area of a rectangle`. His new shed has an area of \(300ft^2\), \(300=?\) and his old shed was \(156ft^2\) less than his new shed, so you can make an equation, \(300=x+156\) OR \(300-156=x\) I prefer \(300=x+156\), so \(300=x+156\) subtract 156 from both sides \(144=x\) So the old storage shed would have been \(144ft^2\).

nightshade0773:

the answer choices it gave me was s+156=300 or s-156=300

Extrinix:

Yeah, so just reverse my equation and replace x for s, \(300=x+156\) Which would be, \(s+156=300\)

nightshade0773:

thank u Extrinix

nightshade0773:

but then it says that i have to solve it

nightshade0773:

see Solve this equation to find how many square feet of storage Mr. Morton's old shed had. it has to be in square feet